# Vision-Based CNN Prediction of Sunspot Numbers from SDO/HMI Images

### v2

@@ 2.2 Solar Image Data from SDO/HMI
For this study, we employed HMI continuum intensity observations at 6173 Å, accessed as publicly available full-disk quicklook JPEG images from the SOHO/SDO reprocessing archive (https://soho.nascom.nasa.gov/data/REPROCESSING/Completed/). This choice provides a uniform, readily reproducible image stream suitable for computer-vision ingestion, while preserving the photospheric contrast patterns associated with sunspot morphology. This wavelength samples the photospheric layer of the Sun and provides high-resolution images of the visible solar disk, where sunspots appear as dark regions due to their lower temperature relative to the surrounding photosphere. Continuum images are therefore particularly well suited for studies that require direct correspondence between sunspot morphology and the sunspot number index.
@@ 2.3 Model Architecture and Training Procedure
Figure 3: Schematic overview of the final CNN regression architecture. The model processes $512\times 512$ three-channel HMI continuum quicklook images through stacked $3\times 3$ convolutional layers with batch normalization and max pooling, followed by Global Average Pooling (512) and a dense regression head (Dense 256 + dropout, Dense 128, Output 1).
@@ 4 Discussion and Conclusions
Future work should focus on three directions. First, a synergistic multimodal model would combine an image encoder (e.g., a CNN/ViT operating on HMI continuum images) with complementary temporal predictors (e.g., recent sunspot-number history and other scalar proxies such as F10.7), using feature fusion (early or late fusion) to capture both instantaneous morphology and short-term temporal context. Second, systematic benchmarking across cycles should be pursued via cross-cycle holdouts (e.g., training on Solar Cycle 24 and testing on Cycle 25, and vice versa) and rolling-origin evaluations, to quantify robustness under domain shifts and changing activity regimes. Third, interpretability tools (Grad-CAM, IG) can be expanded into routine quality-control diagnostics, supporting trust and scientific plausibility checks in automated monitoring pipelines.

### v3

@@ 2.2 Solar Image Data from SDO/HMI
For this study, we employed HMI continuum intensity observations at 6173 Å, accessed as publicly available full-disk quicklook JPEG images from the SOHO/SDO reprocessing archive (https://soho.nascom.nasa.gov/data/REPROCESSING/Completed/). This choice provides a standardized and easily accessible image stream suitable for computer-vision ingestion, while preserving the photospheric contrast patterns associated with sunspot morphology. We note, however, that the upstream pipeline used to generate quicklook products (including calibration parameters and rendering choices) may evolve over time; therefore, our use of quicklooks prioritizes practical reproducibility of data access and formatting rather than guaranteeing strict time-invariant photometric calibration. This wavelength samples the photospheric layer of the Sun and provides high-resolution images of the visible solar disk, where sunspots appear as dark regions due to their lower temperature relative to the surrounding photosphere. Continuum images are therefore particularly well suited for studies that require direct correspondence between sunspot morphology and the sunspot number index.
@@ 2.3 Model Architecture and Training Procedure
This splitting strategy is primarily designed to prevent information leakage in a time-ordered dataset. Nevertheless, a purely chronological split does not guarantee identical activity-level distributions across subsets (e.g., different proportions of rising/declining phases), and a hold-out validation subset drawn from the training period does not fully assess generalization under distribution shift. A more rigorous assessment would require blocked cross-validation, cross-cycle holdouts, or cross-instrument evaluation protocols. We therefore interpret the reported test performance as evidence of strong within-dataset generalization, while acknowledging that broader generalization (e.g., across solar-cycle regimes or instrument domains) remains an important direction for future benchmarking.
@@
Figure 3: Schematic overview of the final CNN regression architecture. The model processes $512\times 512$ three-channel HMI continuum quicklook images through stacked $3\times 3$ convolutional layers with batch normalization (BN) and max pooling, followed by Global Average Pooling (GAP; 512) and a dense regression head (Dense 256 + dropout, Dense 128, Output 1).
@@ 4 Discussion and Conclusions
Future work should focus on three directions. First, a synergistic multimodal model would combine an image encoder (e.g., a CNN/ViT operating on HMI continuum images) with complementary temporal predictors (e.g., recent sunspot-number history and other scalar proxies such as F10.7), using feature fusion (early or late fusion) to capture both instantaneous morphology and short-term temporal context. Second, systematic benchmarking across cycles should be pursued via cross-cycle holdouts (e.g., training on Solar Cycle 24 and testing on Cycle 25, and vice versa) and rolling-origin evaluations, to quantify robustness under domain shifts and changing activity regimes. Extending benchmarks across multiple solar cycles requires not only long sunspot-number records but also consistent full-disk imagery. While SDO/HMI provides homogeneous observations from 2010 onward, earlier cycles would require incorporating pre-SDO image archives, such as SoHO-era observations (e.g., MDI continuum/intensity products) or long-term ground-based full-disk programs. A key challenge is cross-instrument domain shift: differences in spatial resolution, photometric response, compression and rendering pipelines, cadence, and data gaps can strongly affect image statistics. Addressing this would likely require careful harmonization (disk normalization, consistent cropping/rescaling, intensity standardization) and explicit cross-instrument evaluation (train on one instrument, test on another), potentially supported by domain adaptation methods. Third, interpretability tools (Grad-CAM, IG) can be expanded into routine quality-control diagnostics, supporting trust and scientific plausibility checks in automated monitoring pipelines.